OpenStudy (anonymous):

Solve: 4p^2+4p+1=0 Solve for both "P's"

OpenStudy (ash2326):

@vanes5623 do you know the expansion of this \[(a+b)^2?\]

OpenStudy (anonymous):

what do you mean?

OpenStudy (ash2326):

\[(a+b)^2=a^2+b^2+2ab\] do you know this?

OpenStudy (anonymous):

i think so

OpenStudy (ash2326):

4p^2+4p+1=0 Can you simplify the left side, use the hint, I gave you!!

OpenStudy (anonymous):

I can't. I'm so confusedd!

OpenStudy (ash2326):

What's the expansion of this? \[(2p+1)^2??\]

OpenStudy (anonymous):

I dont know..?

OpenStudy (ash2326):

Use the formula \[(a+b)^2=a^2+2ab+b^2\] a=2p and b=1 here

OpenStudy (anonymous):

I'm sorrry. I can't do it.

OpenStudy (ash2326):

Try once, simplify just the right side \[(2p+1)^2=(2p)^2+2\times (2p) \times (1)+1^2\]

OpenStudy (anonymous):

I can't. I'm so confusedd!

OpenStudy (ash2326):

What's the square of 2p?

OpenStudy (anonymous):

im only a freshman in high school. I don't understand this

OpenStudy (anonymous):

can i just have the answer.. please

OpenStudy (anonymous):

lol okay for starters on OS never just ask for an answer. That annoys long-time users lol And Algebra I sucked so I feel your pain either way p = -1/2 @ash2326 explained it perfectly to you, I figured it out by working out his examples! haha It's kind of hard to keep up but if you pay close attention it's not so bad :3

OpenStudy (anonymous):

Don't forget to give Ash a medal for taking the time to help you :)